OpenStudy (anonymous):

A line contains the point (–3, 4). If the slope of the line is -2/3, write the equation of the line using slope-intercept form

OpenStudy (carolina4567):

Slope-intercept form of a line: y = mx + b.

OpenStudy (anonymous):

i already knew that

OpenStudy (carolina4567):

Then all you need is to setup :)

OpenStudy (carolina4567):

Y is the slope , and they gave you that , so y is ?

OpenStudy (anonymous):

-2/3?

OpenStudy (hyuna301):

Use ( y - y1) = m ( x - x1)

OpenStudy (carolina4567):

Good -2/3 = m and x which would be the points + b

OpenStudy (anonymous):

so y=-2/3x+4?

OpenStudy (hyuna301):

No , y = -2/3 x + 2